# Accelerating Clinical Evidence Synthesis with Large Language Models

### v1

@@ Discussion
Clinical evidence forms the bedrock of evidence-based medicine, crucial for enhancing healthcare decisions and guiding the discovery and development of new drugs. It often comes from a systematic review of diverse studies found in the literature, encompassing clinical trials and retrospective analyses of real-world data. Yet, the burgeoning expansion of literature databases presents formidable challenges in efficiently identifying, summarizing, and maintaining the currency of this evidence.
@@
The rapid development of large language models (LLMs) and generative AI technologies has generated considerable interest in their potential applications. However, implementing these models in a manner that is collaborative, transparent, and trustworthy poses significant challenges, especially in critical areas such as medicine [45]. For instance, when utilizing LLMs to summarize multiple studies, the summaries often merely echo the findings verbatim, omit crucial details, and fail to adhere to established best practices [9].
@@
This study introduces a clinical evidence synthesis pipeline enhanced by LLMs, named TrialMind. This pipeline is structured in accordance with established medical systematic review protocols, involving steps such as study searching, screening, data/result extraction, and evidence synthesis. At each stage, human experts have the capability to access, monitor, and modify intermediate outputs. This human oversight helps to eliminate errors and prevents their propagation through subsequent stages. Unlike approaches that solely depend on the internal knowledge of LLMs, TrialMind integrates human expertise through in-context learning and chain-of-thought prompting. Additionally, TrialMind extends external knowledge sources to its outputs through retrieval-augmented generation and leveraging external computational tools to enhance the LLM’s reasoning and analytical capabilities. Comparative evaluations of TrialMind and traditional LLM approaches have demonstrated the advantages of this system design in LLM-driven applications within the medical field.
@@
This study also has several limitations. First, despite incorporating multiple techniques, LLMs may still make errors at any stage. Therefore, human oversight and verification remain crucial when implementing TrialMind in practical settings. Second, the prompts used in TrialMind were developed based on prompt engineering experience, suggesting potential for performance enhancement through advanced prompt optimization or by fine-tuning the underlying LLMs to suit specific tasks better. Third, while TrialMind demonstrated effectiveness in study search, screening, and information extraction, the dataset used was limited in size due to the high costs associated with human labeling. Future research could expand on these findings with larger datasets to further validate the method’s effectiveness. Fourth, the study coverage was restricted to publicly available sources from PubMed Central, which provides structured PDFs and XMLs. Many relevant studies are either not available on PubMed or are in formats that entail OCR algorithms as preprocessing, indicating a need for further engineering to incorporate broader data sources. Fifth, although TrialMind illustrated the potential of using advanced LLMs like GPT-4 to streamline clinical evidence synthesis, developing techniques to adapt the pipeline for use with other LLMs could increase its applicability. Finally, while the use of LLMs like GPT-4 can accelerate study screening and data extraction, the associated costs and processing times may present bottlenecks in some scenarios. Future enhancements that improve efficiency or utilize localized, specialized smaller models could increase practical utility.
@@ Methods
Further, we favored studies that involved a larger number of individual studies, providing a broader base of evidence. However, we excluded network analysis studies and meta-analyses that focused solely on prognostic and predictive effects, as they did not align with our primary research focus. To maintain a balanced representation, we limited our selection to a maximum of three papers per treatment category. This process culminated in a final dataset comprising 25 papers. This curated collection forms the backbone of our analysis, ensuring a concentrated and pertinent selection of high-quality studies directly relevant to our research objectives. The characterstics of the created dataset is in Extended Table LABEL:tab:data_stats.
@@
TrialMind processes research question inputs using the PICO (Population, Intervention, Comparison, Outcome) framework to define the study’s topic and scope. In our experiments, the title of the target review paper served as the general description. Subsequently, we extracted the PICO elements from the paper’s abstract to detail the specific aspects of the research question.
@@
In our literature search experiments, we assessed performance using overall Recall, aiming to evaluate the effectiveness of different methods in identifying all relevant studies from the PubMed database using APIs [50]. For literature screening, we measured efficacy using Recall@20 and Recall@50, which gauge how well the methods can prioritize target studies at the top of the list, thereby facilitating quicker decisions about which studies to include in evidence synthesis. We constructed the ranking candidate set for each review paper by initially retrieving studies through TrialMind, then refining this list by ranking the relevance of these studies to the target review’s PICO elements using OpenAI embeddings. The top 2000 relevant studies were kept. We then ensured all target papers were included in the candidate set to maintain the integrity of our groundtruth data. The final candidate set was then deduplicated to be ranked by the selected methods.
@@
We engaged two groups of annotators for our evaluation: (1) three computer scientists with expertise in AI applications for medicine, and (2) five medical doctors to assess the generated forest plots. Each annotator was asked to evaluate five review studies. For each review, we randomly presented forest plots generated by both the baseline and TrialMind. The annotators were required to determine how closely each generated plot aligned with a reference forest plot taken from the target review paper. Additionally, they were asked to judge which method, the baseline or TrialMind, produced better results in a win/lose assessment. Fig. 5a demonstrates the user interface for this study, which was created with Google Forms.
Extended Fig. 1: The flowchart of the screening process of meta-analyses involved in the MetaSyns dataset.
@@
Table 1: The characteristics of involved meta-analyses papers in the MetaSyns dataset. $N_{1}$: the number of identified studies; $N_{2}$: the number of involved studies after screening; $N_{3}$: the number of involved participants in the studies; P: population; I/C: intervention and comparison; O: outcome measurements.
Ref Topic $N_{1}$ $N_{2}$ $N_{3}$ P I/C O
Valle, et al. 202132 Brachytherapy 2553 150 11322 men treated with a number of salvage treatments for radiorecurrent prostate cancer. Salvage radical prostatectomy (RP), high-intensity focused ultrasound (HIFU), cryotherapy, stereotactic body radiotherapy (SBRT), low–dose-rate (LDR) brachytherapy, and high-dose-rate (HDR) brachytherapy 1. 2-yr and 5-yr Relapse free survival (RFS); 2. severe genitourinary (GU) toxicity; 3. severe gastrointestinal (GI) toxicity
Li et al. 202333 Brachytherapy 11806 15 12773 prostate cancer 1. Brachytherapy (BT); 2. External beam radiotherapy (EBRT); 3. BT+EBRT 1. risk ratios (RRs) of genitourinary (GU) toxicity; 2. RRs of gastrointestinal (GI) toxicity
Hande et al. 202234 Brachytherapy 5499 24 5488 locally advanced cervical cancer 1. Volume-based Brachytherapy (BT); 2. Point-A based BT 1. 3-year disease-free survival (DFS); 2. 3-year local control (LC); 3. 3-year overall survival (OS); 4. severe GU toxicity 5. severe GI toxicity
Wahyuhadi et al. 202244 Cancer Vaccines 326 6 1202 Glioblastoma multiforme (GBM) 1. active immunotherapy: dendritic cell vaccination, peptide vaccination, DNA vaccine, viral vector-based vaccine, antigen non-specific vaccine, autologous tumor cell therapy; 2. standard therapy (surgery, chemotherapy, radiotherapy) 1. Overall survival (OS); 2. Progression-free survival (PFS); 3. post-treatment Karnofsky performance scale (KFS); 4. serious adverse events (AEs); 5. 2-year mortality
Zhou et al. 202045 Checkpoint Inhibitors 2236 30 4971 patients with cancer receiving immune checkpoint inhibitors (ICIs): nivolumab, pembrolizumab, atezolizumab, durvalumab, avelumab, or ipilimumab 1. occurrence of immune-related AEs (irAEs); 2. non-occurrence of irAEs 1. overall survival (OS); 2. progression-free survival (PFS)
Leone et al. 202246 Checkpoint Inhibitors 984 10 5257 advanced esophageal squamous cell carcinoma (ESCC) immune checkpoint inhibitors (ICIs) 1. overall survival benefit; 2. progression-free survival; 3. overall response rate
Song et al. 202047 Checkpoint Inhibitors 220 147 23761 cancer patients 1. anti-PD-1 & anti-PD-L1 inhibitors; 2. anti-CTLA-4 inhibitors ICI-related AEs (Grade 1-5, Grade 3-5)
ABC Group, 202235 Chemotherapy N/A 10 1183 participants with biopsy-proven transitional cell carcinoma with muscle-invasive bladder cancer who had not received neoadjuvant chemotherapy 1. adjuvant cisplatin-based chemotherapy plus local treatment; 2. local treatment; 3. local treatment then adjuvant cisplatin-based chemotherapy on reccurence 1. overall survival; 2. locoregional recurrence-free survival; 3. metastasis-free survival; 4. overall recurrence-free survival
Lacas et al. 202136 Chemotherapy N/A 107 19805 squamous cell Head and Neck Cancer (MACH-NC) Q1: 1. loco-regional treatment (LRT); 2. LRT+CT; Q2: 1. induction CT + radiotherapy; 2. concomitant CT 1. overall survival; 2. event-free survival; 3. loco-regional failure (LRF); 4. distant failure (DF); 5. cancer and non-cancer mortality
Xia et al. 202037 Chemotherapy N/A 9 36480 triple-negative breast cancer (TNBC) 1. neoadjuvant chemotherapy (NACT); 2. adjuvant chemotherapy (ACT) 1. overall survival; 2. disease-free survival (DFS)
Liu et al. 202141 Hormone Therapy 7506 16 67616 high-risk prostate cancer (HRPCa) 1. neoadjuvant hormone therapy (NHT) with radical prostatectomy (RP); 2. RP alone 1. overall survival; 2. biochemical progression-free survival; 3. cancer-specific survival; 4. disease-free survival; 5. risk rates (RRs) of lymph node involvement; 6. RRs of pathological downstaging; 7. RRs of organ-confinement; 8. RRs of positive surgical margins; 9. RRs of seminal vesicle invasion
Piezzo et al. 202042 Hormone Therapy 685 8 4580 patients with HR-positive/HER2-negative advanced or metastatic breast cancer 1. CDK4/6 inhibitors with endocrine therapy (ET); 2. ET only 1. progression-free survival; 2. overall survival; 3. objective response rate (ORR)
Peleg Hasson et al. 202143 Hormone Therapy 113 6 35680 HR-positive/HER2-positive early-stage breast cancer 1. adjuvant endocrine therapy (tamoxifen); 2. aromatase inhibitors disease-free survival
Kim et al. 202140 Hyperthermia 334 4 653 low-risk papillary thyroid microcarcinomas (PTMCs) 1. thermal ablation; 2. surgery 1. new tumor after treatment; 2. lymph node metastasis; 3. complication
Wang et al. 202239 Hyperthermia 3060 9 1215 T1aN0M0 and T1bN0M0 papillary thyroid carcinoma ultrasound-guided thermal ablation (LA, RFA, or MWA) 1. volume reduction rate; 2. overall diseasee progress rate; 3. complication rate
Spiliotis et al. 202138 Hyperthermia 716 15 2169 liver cancer 1. microwave ablation; 2. radiofrequency ablation 1. complete ablation (CA); 2. local tumor prgression (LTP); 3. intrahepatic distant (IDR); 4. complications
Li et al. 202048 Immune System Modulators 682 9 1752 breast cancer receiving chemotherapy 1. PEGylated granulocyte colony-stimulating factor (G-CSF); 2. G-CSF 1. risk ratios (RRs) of grade ¿= 3 / 4 neutropenia; 2. RRs of febrile neutropenia (FN); 3. time to absolute neutrophil count recovery; 4. grade 4 AEs; 5. RRs of skeletal and/or muscle pain
Wu et al. 202249 Monoclonal Antibodies 116 32 958 HER2-positive patients with NSCLC HER2-targeted therapy 1. objective response rate (ORR); 2. disease control rate (DCR); 3. progression-free survival (PFS)
Li et al. 202150 Monoclonal Antibodies 5861 27 15063 patients with solid tumors 1. anti-PD-1/PD-L1 mAbs; 2. anti-PD-1/PD-L1 mAbs + chemotherapy; 3. standard chemotherapy RR of nephrotoxicity
Zhu et al. 202351 Monoclonal Antibodies 2511 169 22492 cancer antibody–drug conjugates (ADCs) 1. RR of grade ¿= 3 AEs; 2. RR of all grade AE
Masetti et al. 202254 Stem Cell Transplant 2141 9 1448 children with pediatric acute myeloid leukemia (AML) in first complete remission (CR1) 1. allogeneic hematopoietic stem cell transplantation (allo-HSCT); 2.chemotherapy alone 1. overall survival; 2. relapse rate; 3. disease-free survival
Zeng et al. 202155 Stem Cell Transplant 499 15 959 adult philadelphia chromosome positive acute lymphoblastic leukemia in post-remission 1. allogeneic hematopoietic stem cell transplantation (allo-HSCT); 2. tyrosine kinase inhibitor (TKI) combined with chemotherapy 1. overall survival; 2. Relapse-free survival (RFS); 3. Odds Ratio of non-relapsed mortality (NRM); 4. Odds Ratio of non-relapsed survival (NRS)
Gagelmann et al. 202156 Stem Cell Transplant 1050 7 680 patients with FLT3-ITD-mutated acute myeloid leukemia (AML) and received allogeneic stem-cell transplation tyrosine kinase inhibitor (TKI) maintenance therapy 1. relapse-free survival; 2. RRs of relapse; 3. overall survival; 4. non-replase mortality; 5. RRs of chronic graft vs. chronic disease (GVHD)
Yang et al. 202152 T-Cell Transfer Therapy 661 23 350 patients with Relapse or Refractory Multiple Myeloma (RRMM) and treated with CAR-T therapy CAR-T therapy 1. Overall response (OR); 2. complete response rate (CRR); 3. MRD negativity within responders; 4. relapse rate; 5. progression-free survival; 6. overall survival; 7. severe CRS (sCRS); 8. Neurologic toxicity (NT)
Shahzad et al. 202353 T-Cell Transfer Therapy 677 13 57 relapsed/refractory acute myeloid leukemia (RR-AML) CAR-T therapy 1. complete remission (CR); 2. overall response rate; 3. incidence of CRS; 4. incidence of ICANs; 5. incidence of graft-versus-host disease (GVHD)

### v2

@@ Discussion
Clinical evidence forms the bedrock of evidence-based medicine, crucial for enhancing healthcare decisions and guiding the discovery and development of new therapies. It often comes from a systematic review of diverse studies found in the literature, encompassing clinical trials and retrospective analyses of real-world data. Yet, the burgeoning expansion of literature databases presents formidable challenges in efficiently identifying, summarizing, and maintaining the currency of this evidence. For instance, a study by the US Agency for Healthcare Research and Quality (AHRQ) found that half of 17 clinical guidelines became outdated within a couple of years [28].
@@
The rapid development of large language models (LLMs) and AI technologies has generated considerable interest in their potential applications in clinical research [29, 30]. However, most of them focused on an individual aspect of the clinical evidence synthesis process, such as literature search [31, 32], citation screening [33, 34, 35], quality assessment [36], or data extraction [37, 38]. In addition, implementing these models in a manner that is collaborative, transparent, and trustworthy poses significant challenges, especially in critical areas such as medicine [39]. For instance, when utilizing LLMs to summarize evidence from multiple studies, the descriptive summaries often usually merely echo the findings verbatim, omit crucial details, and fail to adhere to established best practices [14]. Besides, when given a set of studies that are irrelevant to the research question, LLMs are prone to produce hallucinations and hence cause misleading evidence [40]. This challenge highlights the need for an integrated pipeline, involving the study search and screening stages, to strategically pick the target studies for analysis [41, 42], or enhanced with human-AI collaboration [43].
@@
This study introduces a clinical evidence synthesis pipeline enhanced by LLMs, named TrialMind. This pipeline is structured in accordance with established medical systematic review protocols, involving steps such as study searching, screening, data/result extraction, and evidence synthesis. At each stage, human experts have the capability to access, monitor, and modify intermediate outputs. This human oversight helps to eliminate errors and prevents their propagation through subsequent stages. Unlike approaches that solely depend on the knowledge of LLMs, TrialMind integrates human expertise through in-context learning and chain-of-thought prompting. Additionally, TrialMind extends external knowledge sources to its outputs through retrieval-augmented generation and leveraging external computational tools to enhance the LLM’s reasoning and analytical capabilities. Comparative evaluations of TrialMind and traditional LLM approaches have demonstrated the advantages of this system design in LLM-driven applications within the medical field.
@@
This study also has several limitations. First, despite incorporating multiple techniques, LLMs may still make errors at any stage. Therefore, human oversight and verification remain crucial when implementing TrialMind in practical settings. Second, the prompts used in TrialMind were developed based on prompt engineering experience, suggesting potential for performance enhancement through advanced prompt optimization or by fine-tuning the underlying LLMs to suit specific tasks better. Third, while TrialMind demonstrated effectiveness in study search, screening, and data extraction, the dataset used was limited in size due to the high costs associated with human labeling. Future research could expand on these findings with larger datasets to further validate the method’s effectiveness. Fourth, the study coverage was restricted to publicly available sources from PubMed Central, which provides structured PDFs and XMLs. Many relevant studies are either not available on PubMed or are in formats that entail OCR algorithms as preprocessing, indicating a need for further engineering to incorporate broader data sources. Fifth, although TrialMind illustrated the potential of using advanced LLMs like GPT-4 to streamline clinical evidence synthesis, developing techniques to adapt the pipeline for use with other LLMs could increase its applicability. Finally, while the use of LLMs like GPT-4 can accelerate study screening and data extraction, the associated costs and processing times may present bottlenecks in some scenarios. Future enhancements that improve efficiency or utilize localized, specialized smaller models could increase practical utility.
@@ Methods
Further, we favored studies that involved a larger number of individual studies, providing a broader base of evidence. However, we excluded network analysis studies and meta-analyses that focused solely on prognostic and predictive effects, as they did not align with our primary research focus. To maintain a balanced representation, we limited our selection to a maximum of three papers per treatment category. This process culminated in a final dataset comprising 100 systematic review papers. This curated collection forms the backbone of our analysis, ensuring a concentrated and pertinent selection of high-quality studies directly relevant to our research objectives.
@@
TrialMind processes research question inputs using the PICO (Population, Intervention, Comparison, Outcome) framework to define the study’s research question. In our experiments, the title of the target review paper served as the general description. Subsequently, we extracted the PICO elements from the paper’s abstract to detail the specific aspects of the research question.
@@
In our literature search experiments, we assessed performance using the overall Recall, aiming to evaluate the effectiveness of different methods in identifying all relevant studies from the PubMed database using APIs [47]. For literature screening, we measured efficacy using Recall@20 and Recall@50, which gauge how well the methods can prioritize target studies at the top of the list, thereby facilitating quicker decisions about which studies to include in evidence synthesis. We constructed the ranking candidate set for each review paper by initially retrieving studies through TrialMind, then refining this list by ranking the relevance of these studies to the target review’s PICO elements using OpenAI embeddings. The top 2,000 relevant studies were kept. We then ensured all target papers were included in the candidate set to maintain the integrity of our ground-truth data. The final candidate set was then deduplicated to be ranked by the selected methods.
@@
We engaged two groups of annotators for our evaluation: (1) three computer scientists with expertise in AI applications for medicine, and (2) five medical doctors to assess the generated forest plots. Each annotator was asked to evaluate five review studies. For each review, we randomly presented forest plots generated by both the baseline and TrialMind. The annotators were required to determine how closely each generated plot aligned with a reference forest plot taken from the target review paper. Additionally, they were asked to judge which method, the baseline or TrialMind, produced better results in a win/lose assessment. Extended Fig. 1 demonstrates the user interface for this study, which was created with Google Forms.
Extended Fig. 1: The study design compares the synthesized clinical evidence from the baseline and TrialMind via human evaluation.
Extended Fig. 2: The flowchart of the screening process of meta-analyses involved in the TrialReviewBench dataset.